OpenStudy (anonymous):

Help mee please!! The formula for that area of a triangle is A = 1/2bh. If the base is 2sqrt5 and the height is 4sqrt 7, what is the area of this triangle in radical form?

OpenStudy (anonymous):

you are being asked to multiply \[\frac{1}{2}\times 2\sqrt{5}\times 4\sqrt{7}\]

OpenStudy (anonymous):

the twos cancel, you get \[4\sqrt{35}\]

OpenStudy (anonymous):

that's it ?

OpenStudy (anonymous):

Its just multiplication.

OpenStudy (anonymous):

You could multiply each number individually and you still get the same answer.

OpenStudy (anonymous):

Answer in decimals is 23.6643 If I do\[2\sqrt{5}*4\sqrt{7}*\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

I still get that number. If I do \[4\sqrt{35}\] I still get that number